[clas_members] [Revised Logentry] RC Daily Summary

daria at jlab.org daria at jlab.org
Thu Sep 8 13:55:01 EDT 2022

Logentry Text:
Yesterday was maintenance day, which took all day shift. 

Target group used it to anneal the target and do a quick shim study which concluded that the shims were already in a pretty optima position. They also took a TE polarisation measurement.

Rafo did an FC calibration run.

Dan Carman rebooted the problematic ECAL crate which had gone the night before. 

Beam returned after 5pm and we spent much of the evening optimising the raster, as the beam at first was very unsteady and tripping a lot. The raster size in y was shrunk a little from the previous nominal (3.25 Vpp -> 3.05 Vpp), and the offsets shifted, to avoid the scraping of the cell that we saw the night before. 

Izzy updated HTCC constants.

Overnight, it was smooth sailing. We ran at 7nA, which pushed the peak DC R1 occupancies quite high (closer to 8%) -- to be reviewed.

16902: 100M events (pinned)
16903: 102M events 
16904: 100M events (pinned)
16905: over 130M events, long run before beam-off.

At the moment the beam is off for Hall C beam calibration. Expected to return soon. 

*** Daily RC meeting shortly, at 2pm ***

Your RC


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4039217
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220908/aa0de5a3/attachment.html>

More information about the clas_members mailing list